# Hemlet: A Heterogeneous Compute-in-Memory Chiplet Architecture for Vision Transformers with Group-Level Parallelism thanks: *Equal contribution

### v1

@@ Abstract
Vision Transformers (ViTs) have established new performance benchmarks in vision tasks such as image recognition and object detection. However, these advancements come with significant demands for memory and computational resources, presenting challenges for hardware deployment. Heterogeneous compute-in-memory (CIM) accelerators have emerged as a promising solution for enabling energy-efficient deployment of ViTs. Despite this potential, monolithic CIM-based designs face scalability issues due to the size limitations of a single chip. To address this challenge, emerging chiplet-based techniques offer a more scalable alternative. However, chiplet designs come with their own costs, as they introduce more expensive communication through network-on-package (NoP) compared to network-on-chip (NoC), which can hinder improvements in throughput.
@@
This work introduces Hemlet—a heterogeneous CIM chiplet system designed to accelerate ViT. Hemlet facilitates flexible resource scaling through the integration of heterogeneous analog CIM (ACIM), digital CIM (DCIM), and Intermediate Data Process (IDP) chiplets. To improve throughput while reducing communication overhead, it employs a group-level parallelism (GLP) mapping strategy and system-level dataflow optimization, achieving speedups ranging from 1.44× to 4.07× across various hardware configurations within the chiplet system. Our evaluation results demonstrate that Hemlet can achieve a throughput of 8.68 TOPS with an energy efficiency of 3.86 TOPS/W.
I Introduction
@@
The Vision Transformer (ViT) [vit] has emerged as a foundational model that replaces convolutional neural network (CNN) with Transformer blocks, achieving state-of-the-art performance in computer vision. However, this improvement in performance comes with significantly higher computational demands, primarily driven by massive vector-matrix multiplications (VMMs). In addition to its computational intensity, ViT models are characterized by larger parameter sizes and require substantial frequent memory access to weights during inference. This frequent data transfer limits both throughput and energy efficiency on traditional von Neumann architectures, presenting a challenge commonly referred to as the ”memory wall” bottleneck.
@@
The compute-in-memory (CIM) paradigm [chi2016prime, shafiee2016isaac, AEPE, yu2018neuro] addresses this bottleneck by performing VMMs directly within memory arrays, thereby reducing data movement and enhancing overall efficiency. Under the CIM paradigm, heterogeneous architectures that combine emerging non-volatile memory (eNVM)-based and SRAM-based CIM designs have been widely explored to address the diverse computation and storage demands of Transformer workloads[x-former, liu2023hardsea, laguna2022hardwaresoftware, jain2023heterogeneous]. This heterogeneous design is primarily driven by the intrinsic trade-offs between different memory technologies. On one hand, eNVM devices, such as resistive random access memory (RRAM), offer high storage density and high energy efficiency [shimeng2021review, jiang2023computememory]. However, their limited write endurance and high write overheads pose challenges for handling dynamically generated VMMs within multi-head attention (MHA)[chakraborty2020resistive, zahoor2020resistive, ankit2020panther]. On the other hand, SRAM-based CIM designs provide faster access and better write endurance, but suffer from lower storage density and higher area overhead. To leverage the complementary strengths of these memory technologies, heterogeneous CIM systems typically employ eNVM-based CIM for static VMMs, which utilize fixed pretrained weights. Meanwhile, SRAM-based CIM is used to handle dynamic VMMs that require frequent operand updates during inference. This division of roles allows for efficient support of both static and runtime-generated computations in Transformer inference.
@@
While the heterogeneous design can effectively address the demands of Transformers and maximize performance by statically storing all pretrained weights on-chip, the large parameter sizes of these models present significant challenges. Specifically, storing all pretrained weights in eNVM-based arrays substantially increases chip area, leading to considerable issues related to fabrication cost and yield [chipletcost]. Furthermore, these designs are often optimized for specific workloads (e.g., ViTs, or customized Transformer variants). As model architectures evolve, these tightly coupled designs necessitate complete redesign, verification, and fabrication cycles, resulting in high recurring engineering costs and extended development timelines. To address the challenges posed by monolithic heterogeneous CIM architectures, chiplet-based design emerges as a promising solution. By integrating multiple chiplets within a single package, these systems enhance both modularity and scalability [shao2019simba, caichipletmapping]. Recent studies have started to investigate the potential of CIM chiplet platforms for accelerating DNN workloads [siam, biglittle, sharma2023heterogeneous, dhingra2025atleusacceleratingtransformersedge, cimchiplet].
@@
However, the scalability offered by chiplet-based systems comes at a cost. Chiplets connected via network-on-package (NoP) face greater data transmission overhead compared to network-on-chip (NoC) on a single chip[LVDS, XSRSerDes, caichipletmapping]. Therefore, optimizing the dataflow to reduce transmission overhead is crucial to maintaining performance in chiplet-based systems. Moreover, we observed that the conventional layer-wise mapping method used for ACIM fails to fully utilize the available hardware resources in real implementations, thereby limiting throughput. This presents an opportunity to compensate for the performance degradation caused by chiplet partitioning.
@@
  3. 3.
    System-level dataflow is optimized to minimize inter-chiplet communication. A fine-grained pipeline is utilized for the communication and computation of static VMMs, effectively masking inter-chiplet communication between ACIM and IDP. Furthermore, the DCIM chiplet architecture is co-optimized with tiled dynamic VMMs, eliminating global buffer access across chiplets. These designs significantly enhance throughput in Hemlet’s heterogeneous chiplet architecture.
  4. 4.
@@
The remainder of this paper is organized as follows: Section II introduces the preliminaries of compute-in-memory and chiplet-based architectures. Section III presents the proposed Hemlet architecture, followed by the GLP weight mapping strategy in Section IV. Section V details the system-level dataflow optimization. Some evaluation results are presented in Section VI and a conclusion is drawn in Section VII.
II Preliminary
@@
II-A Compute-in-Memory
Figure 1: Architecture of (a) Analog CIM and (b) Digital CIM
@@ II-A Compute-in-Memory
Analog Compute-in-Memory (ACIM): Conductance-based ACIM macros perform VMMs by leveraging the resistive properties of emerging non-volatile memory (eNVM) devices such as resistive random access memory (RRAM)[chen2019cmos, yao2020fully], phase change memory (PCM)[ambrogio2023analog, kim2019PCM, le202364], and ferroelectric field-effect transistors (FeFET)[yin2018ferroelectric, ni2018fefet]. This approach is highly appealing for energy-efficient computing due to its high-precision analog multiply-and-accumulate (MAC) capabilities and substantial storage density. As illustrated in Fig.1 (a), matrix elements are encoded into the conductance states of the eNVM cells. Depending on the specific characteristics of the memory devices, a single eNVM cell can represent multi-bit weights across multiple conductance levels, accommodating from 1 to 7 bits per cell [ankit2019puma]. Input vectors are converted into different voltages levels and applied to the arrays via wordlines (WLs), with the resulting column currents representing the MAC outputs according to Kirchhoff’s Law. In general, these inputs can also be high precision, depending on the number of voltage levels employed[zhang2022pimca, lee202328, zhang2023macc]. The resulting analog currents are typically converted back to the digital domain through analog-to-digital converters (ADCs) located at the end of the bitlines for further processing. ACIM has been widely used in neural network inference accelerators[chi2016prime, AEPE, shafiee2016isaac, retransformer, ReBert, lu2023rram], as the VMMs dominate in these applications and pre-trained weights can be programmed into the ACIM array prior to inference, remaining unchanged throughout the process. While ACIM is well-suited for weight stationary implementations, it is not ideal for applications requiring frequent weight updates due to high write costs and endurance limitations.
@@
Digital Compute-in-Memory (DCIM): DCIM, built on SRAM-based memory arrays and featuring in-place digital computation, is evolving rapidly due to its flexibility and scalability to advanced technology nodes[fengbin23isscc, dcim25jssc, autodcim, dac24dcim]. Fig. 1 (b) illustrates the basic architecture of a DCIM array. These arrays consist of DCIM columns that share wordlines (WL) and activation lines (AL). Within each DCIM column, n-bit weights stored in SRAM cells are multiplied with 1-bit inputs on AL using local multipliers for each cycle. The resulting $1\times n$-bit products are accumulated along the columns through a digital adder tree. To support multi-bit input processing, outputs from different cycles are accumulated using shift & accumulation (S/A). Compared to ACIM, DCIM offers much lower rewrite overhead due to the SRAM-based arrays and achieves higher throughput by eliminating the need for ADC conversion. However, it exhibits lower storage density and less energy efficiency.
@@
Heterogeneous CIM for Transformers: ACIM is particularly effective for CNNs because the weights used in VMMs are pre-trained, thus eliminating the need for overwriting. However, the scenario shifts with Transformer-based networks, where certain VMMs related to attention are generated on-the-fly, necessitating rewrites in ACIM-only implementations. On the other side, the rapid scaling of operations in Transformers demands high energy efficiency, making DCIM-only accelerators less optimal. Consequently, efficient Transformer inference calls for a heterogeneous integration of ACIM (for static VMMs) and DCIM (for dynamic VMMs) [liu2023hardsea, li2023h3datten]. Fig. 2 illustrates how computations within a Transformer block are mapped onto the heterogeneous CIM system. Specifically, blue blocks denote the workload executed on ACIM, including the linear projections $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, $\mathbf{W_{V}}$, and $\mathbf{W_{O}}$ in multi-head attention (MHA), as well as the two fully connected layers $\mathbf{W_{1}}$ and $\mathbf{W_{2}}$ in the feed-forward network (FFN). Red blocks correspond to dynamic VMMs, namely $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ in MHA, which are handled by DCIM.
Figure 2: Heterogeneous CIM for Transformer Computation
II-B Chiplet architecture
@@
Chiplet architecture is an advanced packaging solution, which integrates multiple dies on an organic substrate or a silicon interposer within a single package[zeppelin, siinterposer, amdchiplet]. Rather than fabricating one large monolithic chip, the system is partitioned into smaller, self‑contained “chiplets” that are interconnected via a network-on-package (NoP). By partitioning a large monolithic die into smaller chiplets, overall manufacturing yield increases dramatically, and the advanced substrates allow total package areas beyond the lithographic reticle limit, enabling larger systems than a single monolithic die. These characteristics directly overcome the size limitations inherent to monolithic chips and align seamlessly with the hardware resource requirements of large‑scale deep neural network (DNN) workloads[shao2019simba, caichipletmapping, biglittle, cimchiplet].
@@
In addition to overcoming monolithic die size limits, chiplet-based designs offer inherent scalability: system capacity can be tuned simply by adding or removing chiplets within the same package. This plug‑and‑play model allows designers to tailor compute, memory, and I/O resources to specific workload demands—ranging from lightweight edge deployments to high‑performance data‑center configurations—without redesigning a large, complex die. As a result, chiplet architectures can grow or shrink in a granular, cost‑effective manner, matching performance targets and market needs with minimal recurring engineering overhead. However, there is still some trade‑offs between chiplet‑based and monolithic accelerator designs [caichipletmapping]. Notably, the communication overhead introduced by NoP causes chiplet architectures to suffer a performance penalty compared to monolithic chips, which in turn has motivated subsequent research on chiplet performance optimization[shao2019simba, caichipletmapping, cimchiplet, biglittle, sharma2023heterogeneous, dhingra2025atleusacceleratingtransformersedge].
Figure 3: Comparison of monolithic and chiplet-based heterogeneous CIM architectures
@@
Figure 4: (a) Overall system architecture integrating ACIM, DCIM, and IDP chiplets interconnected through a network-on-package (NoP); (b) Internal structure of an ACIM chiplet consisting of multiple ACIM Tiles, a local buffer, and NoP router and TX/RX; (c) Tile architecture comprising an array of processing engines (PEs), along with input/output buffers and the adder tree for accumulation, and architecture of a PE with multiple subarrays (SAs), the adder tree, and dedicated buffers; (d) Architecture of a DCIM chiplet comprising multiple DCIM PEs, each equipped with local input/output buffers. The chiplet also integrates a SIMD unit, a control FSM, a chiplet buffer, and a NoP communication module (router and TX/RX) to support inter-chiplet data movement. (e) Design of an Intermediate Data Process (IDP) chiplet equipped with SRAM banks, a SIMD unit, and communication modules.
III Heterogeneous CIM-based Chiplet Architecture
@@
As previously discussed, the conventional monolithic heterogeneous CIM design for ViTs, illustrated in Fig. 3 (a), is constrained by the area limitation of a single chip. A chiplet-based design addresses this issue by assembling multiple smaller chiplets into a larger system. For chiplet-based design, heterogeneous CIM architectures can be realized in two ways. The first design option is intra-chiplet heterogeneity, where different types of CIM macros are integrated within the same chiplet. The system-level capacity is then scaled by homogeneously duplicating this heterogeneous chiplet (Fig. 3 (b)). This design is relatively straightforward. However, a drawback arises from the requirement that different types of CIM macros within one chiplet must be fabricated under the same technology node. As mentioned previously, while DCIM can be scaled to advanced technology nodes (e.g., 7 nm [7nmdcim] and even 3 nm [3nmdcim]), eNVM-based ACIM is still constrained to more mature nodes [H3DTransformer]. This mismatch forces a compromise in process technology, which we refer to as a technology lock-in, preventing the utilization of the more advanced technology node for DCIM. The other design option is inter-chiplet heterogeneity, where each chiplet contains only one type of CIM macros and different types of CIM are placed on separate chiplets and interconnected via NoP (Fig. 3 (c)). This approach circumvents technology lock-in and offers greater design flexibility, allowing each chiplet to adopt the most suitable technology node for its specific requirements.
@@
The detailed Hemlet architecture, illustrated in Fig. 4 (a), comprises three types of specialized chiplets—ACIM, DCIM, and IDP. All chiplets are interconnected through a NoP. To enable seamless inter-chiplet communication, each chiplet integrates a NoP Router along with NoP transmitters (TX) and receivers (RX)—following the design principles of prior works[shao2019simba, siam, biglittle]. In addition, each chiplet integrates a finite state machine (FSM) for control flow, and a single instruction multiple data (SIMD) unit for vector computation (e.g., LayerNorm, GELU, Softmax, accumulation, etc.), and a chiplet buffer for temporary data storage.
Analog CIM (ACIM) Chiplet: The ACIM chiplet, composed of analog RRAM-based ACIM arrays, functions as the computational engine for static VMMs. A sufficient number of ACIM chiplets are allocated to ensure that the weight matrices of the entire ViT model can be fully stored. Each ACIM chiplet features a three-layer architecture comprising Tiles, processing engines (PEs), and subarrays (SAs), as illustrated in Fig. 4 (b)-(c). This hierarchical design is consistent with the approach used in most CIM architectures [shafiee2016isaac, dnn+neurosim, micsim].
Digital CIM (DCIM) Chiplet: DCIM chiplets utilize SRAM-based digital CIM macros as processing engines (PEs) and feature a flattened design, as illustrated in Fig. 4 (d). The DCIM chiplets are shared among the MHA modules from different blocks of the ViT, resulting in lower hardware occupancy compared to the ACIM chiplets. The capacity of the DCIM is tailored to align with the granularity of the operation for dynamic VMMs, which will be discussed in detail in Section V.
@@
IV Group-Level Parallelism (GLP) Weight Mapping Strategy
Figure 5: Motivation for group-level parallelism (GLP). (a) Typical CIM subarray with MUXs and shared ADCs; (b) A group of columns sharing one ADC and MUX is defined as a “Group”; (c) Time-multiplexed column access severely limits throughput, activating only one column per group per cycle; (d) system-wide ADC under-utilization under the layer-wise mapping method.
IV-A Group-Level Parallelism: Motivation
While static VMMs dominate Transformer operations and consequently consume substantial hardware resources in chiplet systems, conventional layer-wise mapping methods—where weights from a single layer are sequentially mapped to the columns of an ACIM array—often do not fully exploit the available ACIM resources during runtime. As described in Section II-A, although VMMs are performed in the analog domain, ADCs are necessary to convert the results back for further processing. However, in practical implementations, a single ADC is typically shared across multiple memory columns due to physical layout constraints—specifically, ADCs are significantly wider than the pitch of the RRAM columns [shimengreview]. To match the column pitch and reduce ADC area overhead, memory columns are grouped and time-multiplexed via a multiplexer (MUX), allowing only one column per group to access the ADC in each cycle (Fig. 5 (b)). While computation occurs concurrently across all columns, the shared ADC digitization process is inherently serialized through multiplexing, significantly constraining the throughput per subarray, shown in Fig. 5 (c).

### v2

@@ Abstract.
Vision Transformers (ViTs) have established new performance benchmarks in vision tasks such as image recognition and object detection. However, these advancements come with significant demands for memory and computational resources, presenting challenges for hardware deployment. Heterogeneous compute-in-memory (CIM) accelerators have emerged as a promising solution for enabling energy-efficient deployment of ViTs. Despite this potential, monolithic CIM-based designs face scalability issues due to the size limitations of a single chip. To address this challenge, emerging chiplet-based techniques offer a more scalable alternative. However, chiplet designs come with their own costs, as they introduce expensive communication, which can hinder improvements in throughput.
@@
This work introduces Hemlet—a heterogeneous CIM chiplet system designed to accelerate ViT. Hemlet facilitates flexible resource scaling through the integration of heterogeneous analog CIM (ACIM), digital CIM (DCIM), and Intermediate Data Process (IDP) chiplets. To improve throughput while reducing communication overhead, it employs a group-level parallelism (GLP) mapping strategy and system-level dataflow optimization, achieving speedups ranging from 1.89× to 4.47× across various hardware configurations within the chiplet system. Our evaluation results demonstrate that Hemlet can achieve a throughput of 9.24 TOPS with an energy efficiency of 4.98 TOPS/W.
Compute-in-memory, Chiplet, Heterogeneous Computing, Mapping
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware Emerging architecturesccs: Computing methodologies Computer visionccs: Computer systems organization Heterogeneous (hybrid) systems
1. Introduction
@@
The Vision Transformer (ViT) (vit) has emerged as a foundational model that replaces convolutional neural network (CNN) with Transformer blocks, achieving state-of-the-art performance in computer vision. However, this improvement in performance comes with significantly higher computational demands, primarily driven by massive vector-matrix multiplications (VMMs). In addition to its computational intensity, ViT models are characterized by larger parameter sizes and require substantial frequent memory access to weights during inference. This frequent data transfer limits both throughput and energy efficiency on traditional von Neumann architectures, presenting a challenge commonly referred to as the ”memory wall” bottleneck.
@@
The compute-in-memory (CIM) paradigm (chi2016prime; shafiee2016isaac; AEPE; yu2018neuro) addresses this bottleneck by performing VMMs directly within memory arrays, thereby reducing data movement and enhancing overall efficiency. Under the CIM paradigm, heterogeneous architectures that combine emerging non-volatile memory (eNVM)-based and SRAM-based CIM designs have been widely explored to address the diverse computation and storage demands of Transformer workloads(x-former; liu2023hardsea; laguna2022hardwaresoftware; jain2023heterogeneous). This heterogeneous design is primarily driven by the intrinsic trade-offs between different memory technologies. On one hand, eNVM devices, such as resistive random access memory (RRAM), offer high storage density and high energy efficiency (shimeng2021review; jiang2023computememory). However, their limited write endurance and high write overheads pose challenges for handling dynamically generated VMMs within multi-head attention (MHA)(chakraborty2020resistive; zahoor2020resistive; ankit2020panther). On the other hand, SRAM-based CIM designs provide faster access and better write endurance, but suffer from lower storage density and higher area overhead. To leverage the complementary strengths of these memory technologies, heterogeneous CIM systems typically employ eNVM-based CIM for static VMMs, which utilize fixed pretrained weights. Meanwhile, SRAM-based CIM is used to handle dynamic VMMs that require frequent operand updates during inference. This division of roles allows for efficient support of both static and runtime-generated computations in Transformer inference.
@@
While the heterogeneous design can effectively address the demands of Transformers and maximize performance by statically storing all pretrained weights on-chip, the large parameter sizes of these models present significant challenges. Specifically, storing all pretrained weights in eNVM-based arrays substantially increases chip area, leading to considerable issues related to fabrication cost and yield (chipletcost). Furthermore, these designs are often optimized for specific workloads (e.g., ViTs, or customized Transformer variants). As model architectures evolve, these tightly coupled designs necessitate complete redesign, verification, and fabrication cycles, resulting in high recurring engineering costs and extended development timelines. To address the challenges posed by monolithic heterogeneous CIM architectures, chiplet-based design emerges as a promising solution. By integrating multiple chiplets within a single package, these systems enhance both modularity and scalability (shao2019simba; caichipletmapping). Recent studies have started to investigate the potential of CIM chiplet platforms for accelerating DNN workloads (siam; biglittle; sharma2023heterogeneous; dhingra2025atleusacceleratingtransformersedge; cimchiplet).
@@
However, the scalability offered by chiplet-based systems comes at a cost. Chiplets connected via network-on-package (NoP) face greater data transmission overhead compared to network-on-chip (NoC) on a single chip(LVDS; XSRSerDes; caichipletmapping). Therefore, optimizing the dataflow to reduce transmission overhead is crucial to maintaining performance in chiplet-based systems. Moreover, we observed that the conventional layer-wise mapping method used for ACIM fails to fully utilize the available hardware resources in real implementations, thereby limiting throughput. This presents an opportunity to compensate for the performance degradation caused by chiplet partitioning.
@@
  3. (3)
    System-level dataflow is optimized to minimize inter-chiplet communication. A fine-grained pipeline is utilized for the communication and computation of static VMMs, effectively masking inter-chiplet communication between ACIM and IDP. Furthermore, the DCIM chiplet architecture is co-optimized with blocked dynamic VMMs, eliminating global buffer access across chiplets. These designs significantly enhance throughput in Hemlet’s heterogeneous chiplet architecture.
  4. (4)
@@
The remainder of this paper is organized as follows: Section 2 introduces the preliminaries of compute-in-memory and chiplet-based architectures. Section 3 presents the proposed Hemlet architecture, followed by the GLP weight mapping strategy in Section 4. Section 5 details the system-level dataflow optimization. Some evaluation results are presented in Section 6 and a conclusion is drawn in Section 7.
2. Preliminary
@@
2.1. Compute-in-Memory
Figure 1. Architecture of (a) Analog CIM and (b) Digital CIM
Figure 2. Heterogeneous CIM for Transformer Computation
@@ 2.1. Compute-in-Memory
Analog Compute-in-Memory (ACIM): Conductance-based ACIM macros perform VMMs by leveraging the resistive properties of emerging non-volatile memory (eNVM) devices such as resistive random access memory (RRAM)(chen2019cmos; yao2020fully), phase change memory (PCM)(ambrogio2023analog; kim2019PCM; le202364), and ferroelectric field-effect transistors (FeFET)(yin2018ferroelectric; ni2018fefet). This approach is highly appealing for energy-efficient computing due to its high-precision analog multiply-and-accumulate (MAC) capabilities and substantial storage density. As illustrated in Fig.1 (a), matrix elements are encoded into the conductance states of the eNVM cells. Depending on the specific characteristics of the memory devices, a single eNVM cell can represent multi-bit weights across multiple conductance levels, accommodating from 1 to 7 bits per cell (ankit2019puma). Input vectors are converted into different voltages levels and applied to the arrays via wordlines (WLs), with the resulting column currents representing the MAC outputs according to Kirchhoff’s Law. In general, these inputs can also be high precision, depending on the number of voltage levels employed(zhang2022pimca; lee202328; zhang2023macc). The resulting analog currents are typically converted back to the digital domain through analog-to-digital converters (ADCs) located at the end of the bitlines for further processing. ACIM has been widely used in neural network inference accelerators(chi2016prime; AEPE; shafiee2016isaac; retransformer; ReBert; lu2023rram), as the VMMs dominate in these applications and pre-trained weights can be programmed into the ACIM array prior to inference, remaining unchanged throughout the process. While ACIM is well-suited for weight stationary implementations, it is not ideal for applications requiring frequent weight updates due to high write costs and endurance limitations.
@@
Digital Compute-in-Memory (DCIM): DCIM, built on SRAM-based memory arrays and featuring in-place digital computation, is evolving rapidly due to its flexibility and scalability to advanced technology nodes(fengbin23isscc; dcim25jssc; autodcim; dac24dcim). Fig. 1 (b) illustrates the basic architecture of a DCIM array. These arrays consist of DCIM columns that share wordlines (WL) and activation lines (AL). Within each DCIM column, n-bit weights stored in SRAM cells are multiplied with 1-bit inputs on AL using local multipliers for each cycle. The resulting $1\times n$-bit products are accumulated along the columns through a digital adder tree. To support multi-bit input processing, outputs from different cycles are accumulated using shift & accumulation (S/A). Compared to ACIM, DCIM offers much lower rewrite overhead due to the SRAM-based arrays and achieves higher throughput by eliminating the need for ADC conversion. However, it exhibits lower storage density and less energy efficiency.
@@
Heterogeneous CIM for Transformers: ACIM is particularly effective for CNNs because the weights used in VMMs are pre-trained, thus eliminating the need for overwriting. However, the scenario shifts with Transformer-based networks, where certain VMMs related to attention are generated on-the-fly, necessitating rewrites in ACIM-only implementations. On the other side, the rapid scaling of operations in Transformers demands high energy efficiency, making DCIM-only accelerators less optimal. Consequently, efficient Transformer inference calls for a heterogeneous integration of ACIM (for static VMMs) and DCIM (for dynamic VMMs) (liu2023hardsea; li2023h3datten). Fig. 2 illustrates how computations within a Transformer block are mapped onto the heterogeneous CIM system. Specifically, blue blocks denote the workload executed on ACIM, including the linear projections $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, $\mathbf{W_{V}}$, and $\mathbf{W_{O}}$ in multi-head attention (MHA), as well as the two fully connected layers $\mathbf{W_{1}}$ and $\mathbf{W_{2}}$ in the feed-forward network (FFN). Red blocks correspond to dynamic VMMs, namely $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ in MHA, which are handled by DCIM.
2.2. Chiplet architecture
@@
Chiplet architecture is an advanced packaging solution, which integrates multiple dies on an organic substrate or a silicon interposer within a single package(zeppelin; siinterposer; amdchiplet). Rather than fabricating one large monolithic chip, the system is partitioned into smaller, self‑contained “chiplets” that are interconnected via a network-on-package (NoP). By partitioning a large monolithic die into smaller chiplets, overall manufacturing yield increases dramatically, and the advanced substrates allow total package areas beyond the lithographic reticle limit, enabling larger systems than a single monolithic die. These characteristics directly overcome the size limitations inherent to monolithic chips and align seamlessly with the hardware resource requirements of large‑scale deep neural network (DNN) workloads(shao2019simba; caichipletmapping; biglittle; cimchiplet).
@@
In addition to overcoming monolithic die size limits, chiplet-based designs offer inherent scalability: system capacity can be tuned simply by adding or removing chiplets within the same package. This plug‑and‑play model allows designers to tailor compute, memory, and I/O resources to specific workload demands—ranging from lightweight edge deployments to high‑performance data‑center configurations—without redesigning a large, complex die. As a result, chiplet architectures can grow or shrink in a granular, cost‑effective manner, matching performance targets and market needs with minimal recurring engineering overhead. However, there is still some trade‑offs between chiplet‑based and monolithic accelerator designs (caichipletmapping). Notably, the communication overhead introduced by NoP causes chiplet architectures to suffer a performance penalty compared to monolithic chips, which in turn has motivated subsequent research on chiplet performance optimization(shao2019simba; caichipletmapping; cimchiplet; biglittle; sharma2023heterogeneous; dhingra2025atleusacceleratingtransformersedge).
Figure 3. Comparison of monolithic and chiplet-based heterogeneous CIM architectures
@@
Figure 4. (a) Overall system architecture integrating ACIM, DCIM, and IDP chiplets interconnected through a network-on-package (NoP); (b) Internal structure of an ACIM chiplet consisting of multiple ACIM PEs, a local buffer, and NoP router and TX/RX; (c) architecture of a processing engine (PE) with multiple subarrays (SAs), the adder tree, and input/output buffers; (d) Architecture of a DCIM chiplet comprising multiple DCIM PEs. The chiplet also integrates a SIMD unit, a control FSM, a chiplet buffer, and a NoP communication module (router and TX/RX) to support inter-chiplet data movement. (e) Design of an Intermediate Data Process (IDP) chiplet equipped with SRAM banks, a SIMD unit, and NoP communication modules.
3. Heterogeneous CIM-based Chiplet Architecture
@@
As previously discussed, the conventional monolithic heterogeneous CIM design for ViTs, illustrated in Fig. 3 (a), is constrained by the area limitation of a single chip. A chiplet-based design addresses this issue by assembling multiple smaller chiplets into a larger system. For chiplet-based design, heterogeneous CIM architectures can be realized in two ways. The first design option is intra-chiplet heterogeneity, where different types of CIM macros are integrated within the same chiplet. The system-level capacity is then scaled by homogeneously duplicating this heterogeneous chiplet (Fig. 3 (b)). This design is relatively straightforward. However, a drawback arises from the requirement that different types of CIM macros within one chiplet must be fabricated under the same technology node. As mentioned previously, while DCIM can be scaled to advanced technology nodes (e.g., 7 nm (7nmdcim) and even 3 nm (3nmdcim)), eNVM-based ACIM is still constrained to more mature nodes (H3DTransformer). This mismatch forces a compromise in process technology, which we refer to as a technology lock-in, preventing the utilization of the more advanced technology node for DCIM. The other design option is inter-chiplet heterogeneity, where each chiplet contains only one type of CIM macros and different types of CIM are placed on separate chiplets and interconnected via NoP (Fig. 3 (c)). This approach circumvents technology lock-in and offers greater design flexibility, allowing each chiplet to adopt the most suitable technology node for its specific requirements.
@@
The detailed Hemlet architecture, illustrated in Fig. 4 (a), comprises three types of specialized chiplets—ACIM, DCIM, and IDP. All chiplets are interconnected through a NoP. To enable seamless inter-chiplet communication, each chiplet integrates a NoP Router along with NoP transmitters (TX) and receivers (RX)—following the design principles of prior works(shao2019simba; siam; biglittle). In addition, each chiplet integrates a finite state machine (FSM) for control flow, and a single instruction multiple data (SIMD) unit for vector computation (e.g., LayerNorm, GELU, Softmax, accumulation, etc.), and a chiplet buffer for temporary data storage.
Analog CIM (ACIM) Chiplet: The ACIM chiplet, composed of analog RRAM-based ACIM arrays, functions as the computational engine for static VMMs. A sufficient number of ACIM chiplets are allocated to ensure that the weight matrices of the entire ViT model can be fully stored. Each ACIM chiplet features a two-layer architecture comprising processing engines (PEs) and subarrays (SAs), as illustrated in Fig. 4 (b)-(c).
Digital CIM (DCIM) Chiplet: DCIM chiplets utilize SRAM-based digital CIM macros as processing engines (PEs), as illustrated in Fig. 4 (d). The DCIM chiplets are shared among the MHA modules from different blocks of the ViT, resulting in lower hardware occupancy compared to the ACIM chiplets. The capacity of the DCIM is tailored to align with the granularity of the operation for dynamic VMMs, which will be discussed in detail in Section 5.
@@
4. Group-Level Parallelism (GLP) Weight Mapping Strategy
Figure 5. Motivation for group-level parallelism (GLP). (a) Typical CIM subarray with MUXs and shared ADCs; (b) A group of columns sharing one ADC and MUX is defined as a “Group”; (c) Time-multiplexed column access severely limits throughput, activating only one column per group per cycle; (d) system-wide ADC under-utilization under the layer-wise mapping method.
4.1. Group-Level Parallelism: Motivation
While static VMMs dominate Transformer operations and consequently consume substantial hardware resources in chiplet systems, conventional layer-wise mapping methods—where weights from a single layer are sequentially mapped to the columns of an ACIM array—often do not fully exploit the available ACIM resources during runtime. As described in Section 2.1, although VMMs are performed in the analog domain, ADCs are necessary to convert the results back for further processing. However, in practical implementations, a single ADC is typically shared across multiple memory columns due to physical layout constraints—specifically, ADCs are significantly wider than the pitch of the RRAM columns (shimengreview). To match the column pitch and reduce ADC area overhead, memory columns are grouped and time-multiplexed via a multiplexer (MUX), allowing only one column per group to access the ADC in each cycle (Fig. 5 (b)). While computation occurs concurrently across all columns, the shared ADC digitization process is inherently serialized through multiplexing, significantly constraining the throughput per subarray, shown in Fig. 5 (c).